OpenStudy (abhisar):

The normality of 10 ml of a 20 V H2O2 solution is ?

OpenStudy (abhisar):

@ParthKohli

OpenStudy (abhisar):

do u know what's 20 V H2O2 ?

Parth (parthkohli):

20 M H2O2? Not sure what 20 V is... Maybe volume/volume percentage?

OpenStudy (abhisar):

ok..wiat !

OpenStudy (abhisar):

so that means that 20 ml of H2O2 in 100 ml of solution..ryt ?

Parth (parthkohli):

If V means %V/V, then sure.

OpenStudy (abhisar):

ok let us suppose it is %v/v ....then how to proceed towards Normality ?

OpenStudy (abhisar):

How to do that without any density given ?

Parth (parthkohli):

Not sure how...

Parth (parthkohli):

Density is required. Without that, not sure at all.

Parth (parthkohli):

What if it is 20 molar H2O2 solution?

OpenStudy (abhisar):

then it means that 20 moles in 1 litre

OpenStudy (abhisar):

20 moles=34*20=680 g

Parth (parthkohli):

Normality = Molarity * n-factor

OpenStudy (abhisar):

yes..what will be the n factor here ?

OpenStudy (abhisar):

@Abmon98 what do u think ?

Parth (parthkohli):

I'm sorry, I haven't really studied this outside acids and bases. I'm Googling to see what the n-factor is in these cases.

OpenStudy (abhisar):

oh i see..it's ok...thanx for the concern buddy !

OpenStudy (abhisar):

btw the answer is 3.58

Parth (parthkohli):

@abb0t @Australopithecus

OpenStudy (abmon98):

Normality of a solution is defined as the number of gram equivalents of the solute dissolved per litre (dm3) of given solution. is V the volume of solution?

Parth (parthkohli):

But the volume is already given as 10 ml, so it has to be something that's not the volume.

OpenStudy (abhisar):

20 volume means 1dm3 of the H2O2 solution will produce 20dm3 of oxygen, where: H2O2 -> O2 + H2: i.e 1:1 stoichiometry. At 298K and 1atm pressure, 1 mole of gas occupies 24.45dm3. So 20/24.45 = 0.818 mol. O2 produced, which means there were 0.818mol H2O2 in 1dm3. Therefore molarity is 0.818M at 298K.

OpenStudy (abhisar):

Source: http://www.thestudentroom.co.uk/showthread.php?t=377055

OpenStudy (abmon98):

check this out @Abhisar i think this well help

OpenStudy (abmon98):

will**

Parth (parthkohli):

And it's an aqueous solution?

OpenStudy (abhisar):

Wow! Thank you very much @Abmon98 !

OpenStudy (abmon98):

your most welcome :D

Parth (parthkohli):

Wow, nice find!

OpenStudy (abhisar):

Thanx @ParthKohli !

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @ParthKohli And it's an aqueous solution? \(\color{blue}{\text{End of Quote}}\) Yes the solution is aqueous